Details.

power on 10:40

ocam no OK.

bad frames. saved TN:20231108_174531

10:56 temperature not yet ready

stop operations. Internet down from 11:15am

12:02 restore activity.

gopt running no disturbane.TN:20231108_192359

@341 x=-11,y=65

Rotation

Bin3

Ok.

Bin4

sloperms ->0

OK

stopAO

2:30PM

-- JuanCarlosGuerra - 08 Nov 2023
Topic revision: r3 - 09 Nov 2023, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback